Solve the equation:
6|3n-10|-5=115

Answer by addingup(3677)

Solve for n:
6|3n-10|-5 = 115 add 5 to both sides and then divide both sides by 6:
|3n-10|= 20
Absolute value, the answer is one of the next two equations:
3n-10 = 20
or
3n-10 = -20
~~~~~~~~~~~~~~~~~~~~
3n = 30
or
3n = -10
~~~~~~~~~~~~~~~~~~~~
n = 10
or
n = -10/3 = -3.333...